A 25ft ladder is put on a wall,20 ft high, how far is the bottom of the ladder away from the wall ?
Answer by mananth(16946)

The ladder ,ground and the wall form a right triangle
ground distance = leg2
height at which ladder touches
Hypotenuse = ladder length
Pythagoras theorem states that

Hypotenuse ^2= leg1^2+leg2^2

Hypotenuse = 25 ft
leg1 = 20 ft

leg2^2= hypotenuse ^2 - leg1^2

Leg2^2= 25 ^2 - 20 ^2
Leg2^2= 625 - 400
Leg2^2= 225
take the square root
leg2= 15 ft

Distance of ladder on ground = 15 ft
m.ananth@hotmail.ca